OpenStudy (anonymous):

a video game system and several games are sold for $652. the cost of the games is 3 times as much as the cost of the system find the cost of the system and the cost of the gmaes

OpenStudy (anonymous):

2,608 because you have to do it 3times so add 652 three times and then add another 652 and then you have you answer

OpenStudy (anonymous):

ok thanks, but i thought I had to divde 652 by 3

OpenStudy (anonymous):

oh wait yes sorry you do have to divide i read the ? wronge

OpenStudy (anonymous):

o so what do i do

OpenStudy (mathlegend):

So, I read it like this... System + Games = $652

OpenStudy (mathlegend):

S + G = 652 G = 652-S

OpenStudy (mathlegend):

Just so I can solve for Games...

OpenStudy (mathlegend):

Since they said that games cost 3 times more than the system...

OpenStudy (mathlegend):

If we distribute what do we get?

OpenStudy (anonymous):

um what step do I do first

OpenStudy (mathlegend):

G = 3(652-G) So we distribute... Meaning what is inside the parenthesis gets multiplied by the 3.

OpenStudy (mathlegend):

So what do you get @Taisha_Casseus ?

OpenStudy (anonymous):

you get 1956 when 3 is multiplied by 652

OpenStudy (mathlegend):

Good..

OpenStudy (mathlegend):

G = 1956 - 3G Now, solve for G.

OpenStudy (anonymous):

1953

OpenStudy (mathlegend):

No... to solve for G we need get the terms with G on one side of the equal sign... and leave the regular number (1956) on one side of the equal side.

OpenStudy (mathlegend):

Lets move that 3G to the other side.... G = 1956 - 3G +3G +3G ---------------- 4G = 1956

OpenStudy (anonymous):

I am sorry I still dont get it

OpenStudy (mathlegend):

This is solving an equation.... we must combine like terms.

OpenStudy (mathlegend):

1G = 1956 - 3G

OpenStudy (mathlegend):

Do you understand why I moved the term with the G over??

OpenStudy (anonymous):

because 1G nd 3G is not like, but i kno thats not why

OpenStudy (mathlegend):

Both have a G attached to them... so they are like terms

OpenStudy (anonymous):

then um no i do not kno can u explain please

OpenStudy (mathlegend):

3b + 4a Are not like terms... because the letter attached to the number is not the same... understand?

OpenStudy (anonymous):

yeah i get that

OpenStudy (mathlegend):

To move something to the other side... you must do the operate operation... the -3G... you see the minus sign is subtraction so we must add to move it to the other side. G = 1956 - 3G +3G +3G ---------------- 4G = 1956

OpenStudy (anonymous):

ok but y is there a minus sign

OpenStudy (mathlegend):

That is from the beginning...

OpenStudy (mathlegend):

S + G = 652 G = 652 - S

OpenStudy (anonymous):

so G=1956 and S=651

OpenStudy (mathlegend):

No... we haven't even solved that equation yet.

OpenStudy (anonymous):

wait so G is not 1956, ok I am really confused right now

OpenStudy (mathlegend):

G = 1956 - 3G +3G +3G ---------------- 4G = 1956 Now you must divide by 4 on both sides... since 4G represents multiplication you do the opposite and divide. 1956/4 = 489 Understand?

OpenStudy (anonymous):

ok i get that you do the opposite of the sign

OpenStudy (mathlegend):

So since we solved the equation... G = $489 Games

OpenStudy (mathlegend):

So now you can divide that by 3... to get your answer for how much 's' the System costs...

OpenStudy (mathlegend):

489 / 3 = 163

OpenStudy (mathlegend):

Game = $489 System = $163

OpenStudy (anonymous):

S=163

OpenStudy (anonymous):

I kinda of get it now thanks, but question is there an easier way to solve the problem

OpenStudy (mathlegend):

Honestly, I do not know of a more simple way.

OpenStudy (anonymous):

ok thanks that really did help